Two trains are headed for each other on a railroad track. One is moving with velocity 100 km/h (east) and the other with velocity 128 km/h (west).
When the trains are 1.2 km apart, the engineers both simultaneously realize the danger of collision and begin braking with an acceleration value of 0.9 m/s^2.
Decide if the trains can stop before collision and if there is no crash, how far apart are the trains?
1
Expert's answer
2020-05-13T09:48:31-0400
"v^2=2(2as)\to \frac{100+128}{3.6}=2\sqrt{(0.9)s}"

"s=1114\ m"

No crash.